Johari pitches a baseball of mass 0.20 kg. The ball arrives at home plate with a speed of 20 m/s and is batted straight back to Johari with a return speed of 15 m/s. What is the magnitude of change in the ball’s momentum?

Answer by ikleyn(53763)

.
Johari pitches a baseball of mass 0.20 kg. The ball arrives at home plate with a speed of 20 m/s
and is batted straight back to Johari with a return speed of 15 m/s.
What is the magnitude of change in the ball’s momentum?
~~~~~~~~~~~~~~~~~~

The momentum before collision with the wall was  mv = 0.2*20 = 4 (kg*m)/s.


The momentum after collision is 0.2*(-15) = -3 (kg*m)/s.

    (the sign is " - ", (minus), because after collision the speed is in opposite direction).


Thus the change of the momentum is 4 (kg*m)/s - (-3 (kg*m)/s) = 7 kg*m/s.


ANSWER.  The change of the momentum is 7 kg*m/s.

Solved.